### v1

@@ 3 $O(\alpha)$-Approximate Non-Adaptive Contracts
   Insert item $(i_{\max},j_{\max})=\arg\max_{(i,j)\in\mathcal{C}}w_{ij}$;
 return policy that executes only this item;
end if
@@
   Sort items in $\mathcal{C}$ by decreasing density $\frac{w_{ij}}{\mu_{ij}}$;
 return policy that executes items in decreasing density order;
end if
Algorithm 1 SKC (Stochastic Knapsack with Costs)
@@
We begin by solving $\Phi(1)$ to obtain an optimal solution $x$ with the properties established in Lemma 3.2: each item $i$ has one choice $j$ such that $x_{ij}=1$, except for possibly one item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*}j}+x_{i^{*}j^{\prime}}=1$. Let $\mathcal{C}$ denote the set of item-choice pairs selected by the LP, and let $r=|\mathcal{C}|$. If there exists an item with two positive choices, we remove these two choices and replace them with a single virtual choice; whenever the algorithm attempts to insert this virtual item, it randomly selects between the two original choices with probabilities proportional to their values in $x$. Since each item now has a single (possibly virtual) choice, we refer to each item-choice pair $(i,j)$ by the item $i$. Note that from the properties of our solution $x$, it holds that $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$ and $\sum_{k=1}^{r}{\mu_{k}}\leq 1$. We then define a threshold $\mathcal{T}(\alpha)$ according to which we decide whether to insert the item with the maximum profit $w_{\text{max}}$. If the threshold is not met, we insert the items in decreasing order of their density $\frac{w_{i}}{\mu_{i}}$, until the knapsack is full or no more items remain in the candidate set $\mathcal{C}$. Our threshold determines whether or not following the greedy ordering is worth the risk of overflow compared to the alternative of inserting the best item. The bigger $w_{\text{max}},\alpha$ are, the more likely we are to insert the best item.
@@
$\mathbb{E}[\textbf{SKC}]\geq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$
@@
First, if $w_{\max}>\mathcal{T}$, the algorithm attempts only item $(i_{\max},j_{\max})$, gaining an expected net reward of $w_{\max}$. Since $w_{\max}>\mathcal{T}$, we have $\mathbb{E}[\textbf{SKC}]=w_{\max}>\mathcal{T}$, as required.
Now, consider the case $w_{\max}\leq\mathcal{T}$. Define $\textbf{SKC}_{k}$ as the contribution of item $k$ to the algorithm’s profit. It holds that
$\operatorname{\mathbb{E}}[\textbf{SKC}]=\sum_{k=1}^{r}{\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]}~.$
We want to lower bound $\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$. If item $k$ is inserted and does not overflow, we obtain a reward of $v_{k}-p_{k}\geq w_{k}$. On the other hand, if the item overflows, we pay a cost of $p_{k}$. By the definition of the IOR parameter (Definition 2.6), $\alpha\geq\frac{p_{k}}{v_{k}\Pr[s_{k}\leq 1]-p_{k}}=p_{k}/w_{k}$, which implies that an overflowing job incurs a cost of $p_{k}\leq\alpha w_{k}\leq\alpha w_{\max}$.
@@
$\operatorname{\mathbb{E}}[\textbf{SKC}_{k}]$ $=(v_{k}-p_{k})\cdot\Pr[k\mbox{ successfully inserted}]-p_{k}\cdot\Pr[k\mbox{ overflows}]$
@@
In the above, this bound is also true for the virtual item $i^{*}$: the item’s expected truncated size is $\mu_{i^{*}}$, and its expected profit conditioned on the event that it was successfully inserted is at least $w_{i^{*}}$.
@@
$\operatorname{\mathbb{E}}[\textbf{SKC}]$ $\geq\sum_{k=1}^{r}{w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)}-\alpha w_{\max}\sum_{k=1}^{r}{\Pr[k\mbox{ overflows}]}$
$\geq\Phi(1)-\sum_{k=1}^{r}{w_{k}}\sum_{j=1}^{k}{\mu_{j}}-\alpha w_{\max}$ $(*)$
$=\Phi(1)-\sum_{1\leq j\leq k\leq r}{w_{k}\mu_{j}}-\alpha w_{\max}$
@@
$\geq\frac{\Phi(1)}{2}-w_{\max}\cdot\left(\sum_{k=1}^{r}{\mu_{k}}+\alpha\right)$
$\geq\frac{\Phi(1)}{2}-w_{\max}(1+\alpha)$ $\sum_{k=1}^{r}{\mu_{k}}\leq 1$
$\geq\frac{\Phi(1)}{4}\geq\frac{\Phi(1)}{4(1+\alpha)}=\mathcal{T}~.$ $w_{\max}\leq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}.$ (4)
In ()(*) we use the fact that $\sum_{k\in[r]}\Pr[k\mbox{ overflows}]\leq 1$ as at most one item overflows, and $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$. In ()(**) we use the decreasing order of density: for every $j<k$ it holds that $w_{k}\mu_{j}\leq w_{j}\mu_{k}$, and thus $w_{k}\mu_{j}\leq\frac{1}{2}\left(w_{k}\mu_{j}+w_{j}\mu_{k}\right)$.
@@
From 3.8 and Lemma 3.7, we conclude:
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{SKC}]}\leq\frac{2\Phi(1)}{\frac{\Phi(1)}{4(1+\alpha)}}=8(1+\alpha)=O(\alpha)~.$
In our proof, we show that our algorithm finds a policy that approximates the LP value up to a factor $O(\alpha)$. In [11], a similar LP helped guide the solution as its value was a proxy (up to constant factors) of the optimal policy for the stochastic knapsack problem. In our case, this LP provides an upper bound to the optimal policy (see Lemma 3.7), but its value can be larger by a factor $\Omega(\alpha)$ than the profit of the optimal policy (see Appendix A). This shows that we cannot hope for a tighter approximation to the LP’s value.
In Section 4, we show that every non-adaptive policy cannot have a better approximation than $O(\alpha)$ to the optimal adaptive policy (Theorem 4.1).
4 Adaptivity Gaps
In this section we study the power of different policies. In Section 4.1, we start by comparing Non-Adaptive policies (Definition 2.3) to Stopping-time Adaptive policies (Definition 2.4). Theorem 4.1 shows that the profit of a principal that uses the latter can be higher by a factor of $\Omega(\alpha)$ (which is tight according to Theorem 3.1). In Section 4.2 we study the power of Fully-Adaptive policies (Definition 2.2), and show they are more powerful than policies that fix the order of jobs and only decide when to stop processing new jobs. Namely, the gap between Fully-Adaptive policies and Stopping-time Adaptive policies can be as high as $\Omega(\alpha^{1/3})$ (Theorem 4.2). We note that in the instances that give the lower bound, all jobs have a single choice, and therefore, in order to incentivize jobs, we need to pay their cost exactly. Specifically, when determining the IOR, we can substitute the expected payment in Definition 2.6 with the jobs’ costs.
4.1 Stopping-time Adaptive vs. Non-Adaptive $\Omega(\alpha)$ Gap
We show that there exists a $\Omega(\alpha)$ gap between the optimal Non-Adaptive policy and the optimal Stopping-time Adaptive policy. Recall that both models require the algorithm to choose an item ordering at the beginning; the stopping-time adaptive policy can then adaptively choose when to halt, while the non-adaptive policy must attempt all items in the chosen order.
@@ 4 Adaptivity Gaps
For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega(\alpha).$
Proof.
For a constant $\alpha$, the claim is trivially true. We prove the claim for every $\alpha>3$. We build an instance $I$ as follows. Consider some $\epsilon<\frac{1}{2}$ and any $\gamma\in(0,\tfrac{1}{2})$, where $\epsilon$ will be used to determine $\alpha$. Our instance has $n$ jobs for $n=\lceil\frac{1}{\epsilon}\rceil$. Each job has value $v=2$ and cost $c=2-\epsilon$.
The size distribution of job $i\in[n]$ is
$s_{i}\;=\;\begin{cases}a_{i},&\text{with probability }1-\epsilon,\\[4.0pt] b_{i},&\text{with probability }\epsilon,\end{cases}$
where
$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
By Definition 2.6, we have
$\alpha=\frac{2-\epsilon}{\epsilon}\;=\;\Theta\!\left(\frac{1}{\epsilon}\right).$
Note that for every $\alpha>3$, we can choose an $\epsilon<1/2$ in order to set the IOR of our instance to be $\alpha$.
Since for every $i$, $\sum_{j=1}^{i-1}a_{j}+b_{i}=1,$ a stopping-time adaptive policy can safely schedule the jobs sequentially until it encounters the first job $i$ whose realized size equals $b_{i}$, and then halt.
We show that since $b_{i}+a_{i^{\prime}}>1$ for every pair $i<i^{\prime}$, attempting to schedule job $i^{\prime}$ without considering the realization of earlier jobs’ size yields negative expected profit. Consequently, the optimal $\mathrm{NON\_ADAPT}$ policy executes at most a single job. Our proof combines these two facts to yield the desired bound.
Formally, consider a stopping‑time adaptive policy that executes jobs in increasing index order ($i=1,2,\ldots$) until it encounters the first job whose size realizes as $s_{i}=b_{i}$, and then stops. Let $X$ denote the number of jobs the policy executes. We have
Since $n\geq 1/\epsilon$, we get that $\operatorname{\mathbb{E}}[X]\geq\frac{1-e^{-1}}{\epsilon}$. Therefore the expected profit is at least
$\mathrm{STOP\_ADAPT}(I)\;=\;\operatorname{\mathbb{E}}[X]\,(v-c)\;\geq\;\frac{1-e^{-1}}{\epsilon}\,\cdot\,\epsilon\;=\;1-e^{-1}.$
Consider now a non‑adaptive policy. This policy fixes the number of jobs to execute, $k$, and an ordering $\pi=i_{1},i_{2},\ldots,i_{k}$. For each $\ell\in[k]$, let $\mathrm{profit}(\ell)$ denote the profit from the $\ell$‑th job in the sequence. Clearly, $\operatorname{\mathbb{E}}[\mathrm{profit}(1)]=\epsilon$ for every non‑empty sequence of jobs. We will show that for every sequence with $k>1$ we have $\operatorname{\mathbb{E}}[\mathrm{profit}(k)]<0$, which implies that it is always better to execute at most a single job, since we can iteratively remove jobs and only improve the expected profit of the policy.
Given a sequence $\pi=i_{1},i_{2},\ldots,i_{k}$ of jobs, consider first the case in which $i_{\ell}>i_{k}$ for some earlier job in the sequence ($\ell<k$). If $i_{k}$ is reached and its size realises as $s_{i_{k}}=b_{i_{k}}$, the job terminates after the deadline, because $s_{i_{\ell}}+s_{i_{k}}\geq a_{i_{\ell}}+b_{i_{k}}>1$. Hence
[…]
$\geq\Pr[i_{k}\text{ is reached}]\Bigl{(}\Pr[s_{i_{k}}=b_{i_{k}}](-c)+\Pr[s_{i_{k}}=a_{i_{k}}](v-c)\Bigr{)}$
$=\Pr[i_{k}\text{ is reached}]\bigl{(}\epsilon(-c)+(1-\epsilon)(v-c)\bigr{)}$
$=-\,\epsilon\,\Pr[i_{k}\text{ is reached}]\;<\;0.$
Now consider the case where $i_{\ell}<i_{k}$ for all $\ell<k$. Let $j=\arg\max_{\ell<k}i_{\ell}$, the largest‑index job among the first k1k-1 scheduled. Note that, $i_{j}<i_{k}$ by our assumption. We have that
$\operatorname{\mathbb{E}}[\mathrm{profit}(k)]$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}]\cdot\Pr[i_{k}\mbox{ is reached}]$
$=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(k)|s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]\cdot\Pr[s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]$
$+\operatorname{\mathbb{E}}[\mathrm{profit}(k)|i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $(v-c)\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$\leq$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-c)\cdot\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-2}(1-\epsilon-2+\epsilon)$
$=$ $-\epsilon\cdot(1-\epsilon)^{k-2}<0.$
In the above, the third equality follows since $i_{k}$ always overflows in case some $s_{i_{\ell}}=b_{i_{\ell}}$ for some $i_{\ell}<i_{k}$, as $b_{i_{\ell}}+a_{i_{k}}>1$. The first inequality follows from the fact that c-c is negative, and $\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\geq\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$ since for the event $s_{i_{j}}=b_{i_{j}}$ and for all other $\ell<k$, $s_{i_{\ell}}=a_{i_{\ell}},$ we have that $i_{k}$ is reached. Since
$\sum_{\ell<k,\ell\neq j}a_{i_{\ell}}\leq\sum_{\ell=1}^{i_{j}-1}a_{\ell}=1-b_{i_{j}},$
and there will be no overflowing job before $i_{k}$ is executed.
We conclude that the optimal non-adaptive policy only executes a single job, and obtains a profit of $\epsilon$, while the optimal stopping-time adaptive policy obtains a profit of 1, as argued above. Therefore,
$\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}\geq\frac{1-e^{-1}}{\epsilon}=\Omega(\alpha).$
4.2 Fully-adaptive vs. Stopping-time Adaptive $\Omega(\alpha^{1/3})$ Gap
We further show that there is an $\Omega(\alpha^{1/3})$ gap between the optimal stopping–time adaptive policy and the optimal fully adaptive policy. Recall that the stopping-time adaptive model requires the algorithm to fix an ordering of the items at the outset, and it halts if all future sequences of jobs yield a negative expected profit.
Theorem 4.2.
For every $\alpha$ there exists an instance $I$ of the knapsack-contracts problem such that
@@
$\frac{\mathrm{ADAPT}(I)}{\mathrm{STOP\_ADAPT}(I)}=\Omega\bigl{(}{\alpha}^{1/3}\bigr{)}.$
@@
In order to prove the above theorem, we define an instance $I$ consisting of an infinite collection of jobs, each with with value $v=15/\epsilon$ and cost $c=15/\epsilon-\epsilon^{2}$. As in Section 4.1, $\epsilon$ is used to determine $\alpha$. Again, as the claim is trivial for bounded $\alpha$, we prove the claim for large values of $\alpha$.
@@
The job set is partitioned into $1/\epsilon$ types of jobs $L_{1},\ldots,L_{1/\epsilon}$, each containing an infinite number of jobs. A job of type $L_{i}$ has the size distribution
@@
where
$a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
for $\gamma<\tfrac{1}{2}$.
By the definition of the IOR parameter (Definition 2.6),
$\alpha=\frac{15/\epsilon-\epsilon^{2}}{\epsilon^{2}}=15/\epsilon^{3}-1=O\left(\frac{1}{\epsilon^{3}}\right)$
@@
We first analyze the expected reward from a fully adaptive policy.
@@
For the instance given above,
$\mathrm{ADAPT}(I)\;\geq\;\frac{1}{2e}.$
@@
Consider an adaptive policy that selects jobs by type $L_{1}$, then type $L_{2}$, and so on until type $L_{1/\epsilon}$, according to the following rule:
  • When selecting jobs of type $L_{i}$, the policy continues to select such jobs as long as $s_{i}=0$.
  • Once $s_{i}=a_{i}$, the policy moves on to select jobs of type $L_{i+1}$ if $i<1/\epsilon$.
@@
Notice that the execution order ensures each executed job finishes within the time budget, so the policy obtains a reward of exactly $v-c=\epsilon^{2}$ for every executed job. Let $X$ be a random variable for the number of jobs the adaptive policy completes. Since the profit from each job is $\epsilon^{2}$, the policy gets a reward of $\operatorname{\mathbb{E}}[X]\cdot\epsilon^{2}$. Let $X_{i}$ be the number of jobs of type $L_{i}$ completed by the policy. Notice that
@@
By the description of the policy,
$\operatorname{\mathbb{E}}\bigl{[}X_{i}\mid\varphi\text{ executes jobs of type }L_{i}\bigr{]}=\frac{1}{\epsilon+\epsilon^{2}}.$
@@
The second inequality follows because in order for the policy to reach jobs of type $L_{1/\epsilon}$, each time a job of type $L_{i}$ (for $i<1/\epsilon$) has its size realized at a value larger than 0, the policy moves on to jobs of type $L_{i+1}$ only if the job’s size is realized at $a_{i}$, which happens with probability $\frac{\epsilon}{\epsilon+\epsilon^{2}}$. Since this must occur at most $1/\epsilon$ times, the probability is at least
$\left(\frac{\epsilon}{\epsilon+\epsilon^{2}}\right)^{1/\epsilon}.$
Therefore,
$\operatorname{\mathbb{E}}[X_{i}]\;\geq\;\frac{1}{e(\epsilon+\epsilon^{2})}\;>\;\frac{1}{2e\epsilon},$
which implies
$\operatorname{\mathbb{E}}[X]=\sum_{i=1}^{1/\epsilon}\operatorname{\mathbb{E}}[X_{i}]\;>\;\frac{1}{2e\epsilon}\cdot\frac{1}{\epsilon}=\frac{1}{2e\epsilon^{2}}.$
This yields
$\mathrm{ADAPT}(I)\;>\;\epsilon^{2}\cdot\frac{1}{2e\epsilon^{2}}\;=\;\frac{1}{2e},$
as desired. ∎
We next bound the expected reward from a stopping-time adaptive policy.
@@
For the instance given above,
$\mathrm{STOP\_ADAPT}(I)\;\leq\;4\epsilon.$

### v2

@@ 3.2 Upper Bound
   Insert item $(i_{\max},j_{\max})=\arg\max_{(i,j)\in\mathcal{C}}w_{ij}$ return policy that executes only this item
@@
 return policy that executes items in $\mathcal{C}$ in decreasing order of density $\frac{w_{ij}}{\mu_{ij}}$
Algorithm 1 NA-SKC (Non-Adaptive Stochastic Knapsack with Costs)
@@
We begin by solving $\Phi(1)$ to obtain an optimal solution $x$ with the properties established in Lemma 3.3: each item $i$ has one choice $j$ such that $x_{ij}=1$, except for possibly one item $i^{*}$ with two choices $j,j^{\prime}$ such that $x_{i^{*}j}+x_{i^{*}j^{\prime}}=1$. Let $\mathcal{C}$ denote the set of item-choice pairs selected by the LP, and let $r=|\mathcal{C}|$. If there exists an item with two positive choices, we remove these two choices and replace them with a single virtual choice; whenever the algorithm attempts to insert this virtual item, it randomly selects between the two original choices with probabilities proportional to their values in $x$. Since each item now has a single (possibly virtual) choice, we refer to each item-choice pair $(i,j)$ by the item $i$. Note that from the properties of our solution $x$, it holds that $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$ and $\sum_{k=1}^{r}{\mu_{k}}\leq 1$. We then define a threshold $\mathcal{T}(\alpha)$ according to which we decide whether to insert the item with the maximum utility $w_{\text{max}}$. If the threshold is not met, we insert the items in decreasing order of their density $\frac{w_{i}}{\mu_{i}}$, until the knapsack is full or no more items remain in the candidate set $\mathcal{C}$. Our threshold determines whether or not following the greedy ordering is worth the risk of overflow compared to the alternative of inserting the best item. The bigger $w_{\text{max}},\alpha$ are, the more likely we are to insert the best item.
@@
$\mathbb{E}[\textbf{NA-SKC}]\geq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$
@@
First, if $w_{\max}>\mathcal{T}$, the algorithm attempts only item $(i_{\max},j_{\max})$, gaining an expected net utility of $w_{\max}$. Since $w_{\max}>\mathcal{T}$, we have $\mathbb{E}[\textbf{NA-SKC}]=w_{\max}>\mathcal{T}$, as required.
Now, consider the case $w_{\max}\leq\mathcal{T}$. Define $\textbf{NA-SKC}_{k}$ as the contribution of item $k$ to the algorithm’s utility. It holds that $\operatorname{\mathbb{E}}[\textbf{NA-SKC}]=\sum_{k=1}^{r}{\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]}.$ We want to lower bound $\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]$. If item $k$ is inserted and does not overflow, we obtain a utility of $v_{k}-p_{k}\geq w_{k}$. On the other hand, if the item overflows, we pay a cost of $p_{k}$. By the definition of the IOR parameter (Definition 2.1), $\alpha\geq\frac{p_{k}}{v_{k}\Pr[s_{k}\leq 1]-p_{k}}=p_{k}/w_{k}$, which implies that an overflowing item incurs a cost of $p_{k}\leq\alpha w_{k}\leq\alpha w_{\max}$.
@@
$\operatorname{\mathbb{E}}[\textbf{NA-SKC}_{k}]$ $=(v_{k}-p_{k})\cdot\Pr[k\mbox{ successfully inserted}]-p_{k}\cdot\Pr[k\mbox{ overflows}]$
@@
In the above, this bound is also true for the virtual item $i^{*}$: the item’s expected truncated size is $\mu_{i^{*}}$, and its expected utility conditioned on the event that it was successfully inserted is at least $w_{i^{*}}$.
@@
$\operatorname{\mathbb{E}}[\textbf{NA-SKC}]$ $\geq\sum_{k=1}^{r}{w_{k}\left(1-\sum_{j=1}^{k}{\mu_{j}}\right)}-\alpha w_{\max}\sum_{k=1}^{r}{\Pr[k\mbox{ overflows}]}$
$\geq\Phi(1)-\sum_{1\leq j\leq k\leq r}{w_{k}\mu_{j}}-\alpha w_{\max}$ $(*)$
@@
$\geq\frac{\Phi(1)}{2}-w_{\max}(1+\alpha)\geq\frac{\Phi(1)}{4}\geq\mathcal{T}~,$ (3)
In ()(*) we use the fact that $\sum_{k\in[r]}\Pr[k\mbox{ overflows}]\leq 1$ as at most one item overflows, and $\sum_{k=1}^{r}{w_{k}}=\Phi(1)$. In ()(**) we use the decreasing order of density: for every $j<k$ it holds that $w_{k}\mu_{j}\leq w_{j}\mu_{k}$, and thus $w_{k}\mu_{j}\leq\frac{1}{2}\left(w_{k}\mu_{j}+w_{j}\mu_{k}\right)$. In the final two inequalities we use the fact that $\sum_{k=1}^{r}\mu_{k}\leq 1$ and $w_{\max}\leq\mathcal{T}=\frac{\Phi(1)}{4(1+\alpha)}$.
@@
From 3.9 and lemma 3.8, we conclude:
$\frac{\mathrm{ADAPT}}{\operatorname{\mathbb{E}}[\textbf{NA-SKC}]}\leq\frac{2\Phi(1)}{\frac{\Phi(1)}{4(1+\alpha)}}=8(1+\alpha)=O(\alpha)~.$
In our proof, we show that our algorithm finds a policy that approximates the LP value up to a factor $O(\alpha)$. In [10], a similar LP helped guide the solution as its value was a proxy (up to constant factors) of the optimal policy for the stochastic knapsack problem. In our case, this LP provides an upper bound to the optimal policy (see Lemma 3.8), but its value can be larger by a factor $\Omega(\alpha)$ than the utility of the optimal policy (see Section D.3). This shows that we cannot hope for a tighter approximation to the LP’s value.
4 A Constant Approximation with Limited Adaptivity
In this section, we study the power of limited adaptivity, namely skipping-adaptive and stopping-adaptive policies. In Section 4.1, we prove an $\Omega(\alpha^{1/3})$ gap between stopping-adaptive and skipping-adaptive policies. In Section 4.2, we present an $O(1)$-approximation skipping-adaptive algorithm, highlighting the power of adaptively skipping items in a given ordering.
4.1 Lower Bound for Stopping-Adaptive Policies
Recall that the stopping-adaptive model requires the algorithm to fix a non-adaptive ordering of the items at the outset, and it halts if all future sequences of items yield a negative expected utility. This is arguably the simplest mildly-adaptive class of policies. The following theorem shows that allowing the algorithm to adaptively halt is not sufficient to achieve a constant approximation to the optimal fully-adaptive policy in our model. In fact, the gap already exists between stopping-adaptive and skipping-adaptive policies.
@@ 4.1 Lower Bound for Stopping-Adaptive Policies
For every $\alpha$ there exists an instance $I$ of the MSKC problem such that
@@
$\frac{\mathrm{SKIP\_ADAPT}(I)}{\mathrm{STOP\_ADAPT}(I)}=\Omega\bigl({\alpha}^{1/3}\bigr).$
@@
In order to prove the above theorem, we define an instance $I$ consisting of an infinite collection of items, each with value $v=15/\epsilon$ and cost $p=15/\epsilon-\epsilon^{2}$. As in Section 3.1, $\epsilon$ is used to determine $\alpha$. Again, as the claim is trivial for bounded $\alpha$, we prove the claim for large values of $\alpha$.
@@
The item set is partitioned into $1/\epsilon$ types of items $L_{1},\ldots,L_{1/\epsilon}$, each containing an infinite number of items. An item of type $L_{i}$ has the size distribution
@@
where $a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}\qquad$ for $\gamma<\tfrac{1}{2}$.
By the definition of the IOR parameter (definition 2.1), $\alpha=\frac{15/\epsilon-\epsilon^{2}}{\epsilon^{2}}=15/\epsilon^{3}-1=O\left(\frac{1}{\epsilon^{3}}\right)$.
@@
We first analyze the expected utility from a fully-adaptive policy.
@@
For the instance given above, $\mathrm{SKIP\_ADAPT}(I)\;\geq\;\frac{1}{2e}.$
@@
Consider a skipping-policy that selects items by type $L_{1}$, then type $L_{2}$, and so on until type $L_{1/\epsilon}$, according to the following rule:
  • When selecting items of type $L_{i}$, the policy continues to select such items as long as $s_{i}=0$.
  • Once $s_{i}=a_{i}$, the policy moves on to select items of type $L_{i+1}$ if $i<1/\epsilon$.
@@
Notice that the execution order ensures each inserted item finishes and does not overflow, so the policy obtains a utility of exactly $v-p=\epsilon^{2}$ for every inserted item. Let $X$ be a random variable for the number of items the adaptive policy successfully completes. Since the utility from each item is $\epsilon^{2}$, the policy gets a utility of $\operatorname{\mathbb{E}}[X]\cdot\epsilon^{2}$. Let $X_{i}$ be the number of items of type $L_{i}$ completed by the policy. Notice that
@@
By the description of the policy, $\operatorname{\mathbb{E}}\bigl[X_{i}\mid\varphi\text{ inserts items of type }L_{i}\bigr]=\frac{1}{\epsilon+\epsilon^{2}}.$
@@
The second inequality follows because in order for the policy to reach items of type $L_{1/\epsilon}$, each time an item of type $L_{i}$ (for $i<1/\epsilon$) has its size realized at a value larger than 0, the policy moves on to items of type $L_{i+1}$ only if the item’s size is realized at $a_{i}$, which happens with probability $\frac{\epsilon}{\epsilon+\epsilon^{2}}$. Since this must occur at most $1/\epsilon$ times, the probability is at least $\left(\frac{\epsilon}{\epsilon+\epsilon^{2}}\right)^{1/\epsilon}.$ Therefore, $\operatorname{\mathbb{E}}[X_{i}]\;\geq\;\frac{1}{e(\epsilon+\epsilon^{2})}\;>\;\frac{1}{2e\epsilon},$ which implies $\operatorname{\mathbb{E}}[X]=\sum_{i=1}^{1/\epsilon}\operatorname{\mathbb{E}}[X_{i}]\;>\;\frac{1}{2e\epsilon}\cdot\frac{1}{\epsilon}=\frac{1}{2e\epsilon^{2}}.$ This yields $\mathrm{ADAPT}(I)\;>\;\epsilon^{2}\cdot\frac{1}{2e\epsilon^{2}}\;=\;\frac{1}{2e},$ as desired. ∎
We next bound the expected utility from a stopping-adaptive policy.
@@
For the instance given above, $\mathrm{STOP\_ADAPT}(I)\;\leq\;4\epsilon.$